# Bit Threads: From Entanglement to Geometric Entropies

### v1

@@ 3 Extensions to other entropies
Bit threads can be related to other notions of entropy. In this section we specifically attempt to relate it to black hole (BH) entropy and differential entropy. Since the CPS formalism is commonly applied in the context of black hole thermodynamics [16, 15], a natural extension is to derive a flow representation for black hole entropy. The same thing can be done for differential entropy [28, 29], which is defined in both holographic and non-holographic spacetimes.
@@ 3.1 Black hole entropy
The entropy of a black hole has been a topic of interest for a long time since Bekenstein [30] to the present day. Among many different approaches, the approach to define the black hole entropy as the Noether charge using the CPS formalism developed by Wald [15, 16] is a well established and standard notion of entropy, known as the ‘Wald entropy’. We will construct a flow corresponding to the Wald entropy of a black hole. Following the procedure described in section 2.2, we can find a divergenceless vector field. The only difference will be the norm bound at the RT surface. As we are interested in the black hole entropy, the norm bound should be on a minimal/extremal surface homologous to the black hole horizon (or, more specifically the bifurcation surface). So the relevant condition becomes $v^{a}=-n_{\mathcal{H}}^{a}/4G_{N}$ at the black hole horizon, $n_{\mathcal{H}}^{a}$ being the unit normal to the horizon. Similar to the AdS-Rindler case, the minus sign emphasises that the direction of $v^{a}$ is towards the horizon [^8] If we work with a Poincaré like coordinate system, then $v^{a}$ is naturally towards the horizon and we do not need to put the minus sign.. Then, the BH entropy is given by,
@@
Here, ‘$M$’ and ‘$J$’ are the mass and angular momentum of the BTZ black hole respectively; $l$ is the AdS radius. For simplicity, we will consider a non-rotating BTZ black hole, i.e., $J=0$ and $\mathcal{N}^{\phi}=0$. Now, the metric becomes,
@@
where, $f(r)=1-\dfrac{2M}{r}+r^{2}$ with AdS radius $L=1$. Considering spherical symmetries we can suppress the last term in (3.1). Now the metric becomes,
@@
The rest of the procedure is similar to the last case. Like the previous one, Schwarzschild-AdS is also a solution of the Einstein-Hilbert gravity. Then we can use (2.20) to have the codim-2 current. Parametric variation gives us the different components of the variations of the metric as follows,
@@
This $d$-form or the corresponding vector $v^{a}$ related to it via (2.13) gives the black hole entropy upon integration on a codimension-2 surface homologous to the horizon (or more specifically the bifurcation point).
We can think of these flows as the conserved current corresponding to different gravitational charges, and the divergenceless condition is the condition for conservation of some charges similar to the continuity equation.
3.2 Entropy of cosmological horizon
Cosmological horizons, like those in de Sitter space, represent another important class of causal boundaries in spacetime with associated thermodynamic properties. Similar to black hole horizons, cosmological horizons emit Hawking radiation with a characteristic temperature and possess an entropy proportional to their area. Here, we explore how the bit thread formalism can be extended to describe the entropy of cosmological horizons. The de Sitter spacetime in static coordinates is described by the metric:
$ds^{2}=-\left(\mathfrak{a}^{2}-\frac{r^{2}}{L_{dS}^{2}}\right)dt^{2}+\left(\mathfrak{a}^{2}-\frac{r^{2}}{L_{dS}^{2}}\right)^{-1}dr^{2}+r^{2}d\Omega^{2}_{d-2},$ (3.8)
where, $\mathfrak{a}$ is some parameter, $L_{dS}$ is the de Sitter radius related to the cosmological constant by $\Lambda=\frac{(d-1)(d-2)}{2L_{dS}^{2}}$. The cosmological horizon is located at $r=\mathfrak{a}L_{dS}$, and its entropy is given by the Gibbons-Hawking formula:
$S_{dS}=\frac{{\cal{A}}(horizon)}{4G_{N}}=\frac{\Omega_{d-2}L_{dS}^{d-2}}{4G_{N}},$ (3.9)
where $\Omega_{d-2}$ is the volume of a unit (d2)(d-2)-sphere.
To construct a bit thread representation of this entropy, we follow the same approach used for black hole horizons. We use the timelike Killing vector $\xi^{a}=(\partial_{t})^{a}$ and apply the covariant phase space formalism to obtain the codimension-2 form $\mathbf{k}_{\xi}$. For parametric variations of the metric with respect to the de Sitter radius $L_{dS}$, we can find:
$k_{\xi}^{tr}=\frac{1}{8\pi G_{N}}\dfrac{\mathfrak{a}\hat{\delta}\mathfrak{a}}{r},$
which leads to the current form:
$j_{\xi}^{tr}=\frac{1}{16\pi G_{N}}\dfrac{\mathfrak{a}^{2}}{r}.$ (3.10)
Using the mapping in equation (2.13), we can construct a divergenceless vector field:
$v^{a}=\left(-v^{r},v^{\Omega}\right)=-\dfrac{1}{4G_{N}}\left(\frac{\mathfrak{a}L_{dS}}{r}\sqrt{\mathfrak{a}^{2}-\frac{r^{2}}{L_{dS}^{2}}},0\right).$
In the standard dS static patch metric we generally take $\mathfrak{a}=1$ and it can be obtained by rescaling the coordinates in (3.8). Then the above vector field becomes
$v^{a}=-\dfrac{1}{4G_{N}}\left(\frac{L_{dS}}{r}\sqrt{1-\frac{r^{2}}{L_{dS}^{2}}},0\right)$ (3.11)
This vector field satisfies the necessary conditions for a valid bit thread configuration: it is divergenceless, and proportional to the unit normal $n_{\cal{C}}^{a}$ on the cosmological horizon. When integrated over the horizon, it correctly reproduces the Gibbons-Hawking entropy or Wald entropy in more general cases.
An interesting feature of cosmological horizons is that they are observer-dependent. Different observers in de Sitter space perceive different cosmological horizons. This observer dependence is reflected in the bit thread configuration, which is tied to the specific static patch associated with a particular observer.
@@ 3.3 Differential entropy
In the AdS/CFT correspondence, entanglement entropy in the boundary theory is computed via the Ryu-Takayanagi prescription, which asserts that the entanglement entropy of a boundary region is equal to the minimum area of a surface in the bulk spacetime. This gives us important insights into how quantum information is encoded in spacetime itself. From the gravitational point of view, entanglement entropy generalises from individual intervals to overlapping boundary intervals. Each of them has an entanglement entropy given by the Ryu-Takayanagi formula. Taking the continuum limit of differences between these entropies, one introduces differential entropy [29]:
$S_{diff}=\int d\lambda\frac{\partial S\big{(}\gamma_{L}(\lambda),\gamma_{R}(\lambda)\big{)}}{\partial\lambda}.$
For particular circumstances, this integral provides the gravitational entropy of a bulk “hole”. Namely, differential entropy happens to be equal to the area of the hole divided by $4G_{N}$,
@@
Here, we show that the differential entropy corresponding to a hole in a spacetime can be obtained using the flow prescription. The vector field is derived from CPS formalism with constraints $|v|\leq 1/4G_{N}$ and $|v^{a}n_{a}|=1$ at the boundary of the hole. This $v^{a}$ is divergenceless by definition. So, we can rewrite the differential entropy in terms of the threads as
@@
where $\mathfrak{b}$ is the boundary of the hole and $v=\sqrt{\mathfrak{h_{b}}}\,|n_{a}v^{a}|$. We will look into the explicit construction of the vector field in spherical Rindler space.
@@
The concept of spherical Rindler space generalises the foundational framework of conventional Rindler space that describes a family of observers undergoing radial acceleration away from a shared central point in Minkowski spacetime, establishing causal isolation from a spherical region of radius $R_{0}$ [28]. Unlike linear Rindler acceleration which produces planar horizons, this spherical generalisation creates an observationally inaccessible domain bounded by a perfectly symmetrical horizon surface. The spherical Rindler horizon manifests as a topological boundary separating the causally connected exterior region from the interior zone that remains perpetually hidden to all accelerated observers. This horizon’s geometrical properties maintain exact spherical symmetry, with its characteristic scale being precisely determined by the parameter $R_{0}$ that defines the radius of the inaccessible central region. The $(d+1)$-dimensional metric is given by [28]:
@@
where, $d\Omega_{d-1}^{2}$ denotes the metric on a unit (d1)(d-1)-sphere. This metric structure reveals the intrinsic relationship between the acceleration parameter $R_{0}$, the radial coordinate $r$, and the temporal component that governs the proper time experience of the spherical Rindler observers. The coordinate singularity at $r=R_{0}$ mathematically encodes the causal boundary beyond which no information can propagate to the accelerating observers, establishing this radius both as geometrical and informational horizon. The authors of [28] showed that this horizon gives rise to a notion of entropy proportional to the area of the horizon, which was named as ‘differential entropy’. To obtain the entropy associated with this hole, we generally take the near horizon limit $r\rightarrow 0$. The metric approaches in the limit
@@
Now we can find a constant time slice in this geometry and do the Iyer-Wald formalism to find the entropy of the hole which is essentially given by (3.12). We can also find a flow or bit thread configuration by proceeding in the way discussed in section 2.2. In the near horizon geometry we take the radius of the hole $R_{0}$ as the parameter to be used for the parametric variation of the metric. Using (2.20) and (2.7), we find ,
@@
for the near horizon geometry of the spherical Rindler space. Here $\mathfrak{c}$ is a function of $R_{0}$. Now using the relations (2.7) and (2.13), we find the vector field,
@@
Notice that the vector field is constant along the $r$ direction and the minus sign tells that it ends on the horizon of the hole. This reproduces the entropy of the hole as given by (3.12). So we can see that it is possible to construct a flow based picture for differential entropy. Moreover, from the above construction we can say that the CPS formalism can also be applied to obtain the entropy of hole in a spacetime.
@@
Spherical AdS-Rindler space can be viewed as a spherical Rindler space embedded in an asymptotically AdS spacetime. This space is holographically dual to the ultraviolet part of the boundary field theory, described by a finite time strip. [29]. This enables us to study a hole in a spacetime holographically. Analogous to the spherical Rindler space, this spacetime also has a family of accelerated observers whose worldlines are causally disconnected from a spherical hole in AdS. The metric of such spacetime is given by [29],
$ds^{2}=l_{AdS}^{2}\left(-\sinh^{2}{r}dt^{2}+dr^{2}\right)+\left(R_{0}\cosh{r}+\sqrt{R_{0}^{2}+l_{AdS}^{2}}\sinh{r}\cosh{t}\right)^{2}d\Omega_{d-1}^{2}\,$, (3.18)
where $R_{0}$ is the coordinate radius of the hole. For convenience, we can take the AdS radius to be $1$. The horizon is at $r=0$ similar to the spherical Rindler case. Taking the near horizon limit $r\rightarrow 0$, we end up with same metric (3.15):
@@
in the leading order. So, in this geometry the flow will be same as given in (3.17). This flow reproduces the gravitational entropy of the hole, viz. ‘differential’ entropy. This method is only applicable in the near horizon limit of the hole. If one wants to construct a thread configuration, one needs to use the notion of covariant bit threads as formulated in [4].
@@ 4 First laws of entropy using bit-threads
The CPS formalism helps us to construct a representative flow for a valid bit thread configuration. So one can wonder if it is possible to represent different thermodynamical laws in terms of the thread picture. In this section, we will mainly focus on the first law of thermodynamics for the above mentioned entropies. Before proceeding further, recall that the symplectic current vanishes for Killing symmetries, i.e., $\boldsymbol{\omega}=0=\mathrm{d}\mathbf{k}_{\eta}$. So, we can construct a vector field from the codim-2 current as
@@
which is similar to the map (2.13), and ensures that $v^{a}_{th}$ is divergenceless. We can also impose the norm bound as we did for $v^{a}$ in the previous section. Then we end up with a ‘thermodynamic variation flow’. This flow when integrated at the boundary of a Cauchy surface, results in variations of different thermodynamic quantities.
@@
For a stationary, rotating, charge neutral, axially symmetric black hole with non-degenerate Killing vectors and a bifurcation surface, we derive the first law of the black hole thermodynamics using the covariant phase space formalism as briefly discussed in appendix B.3. For this kind of black holes, the Killing vectors associated with the horizon, time translation and rotation are related via the equation (B.21),
@@
where, $\Omega_{H}$ represents the angular velocity of the horizon. Now, it is easy to find out the relation between the codim-2 currents corresponding to the Killing vectors and is given by,
@@
Here we have used the notation (A.2). For simplicity we will omit the suffix ‘$th$’ hereafter. These fields also follow a similar relation like (4.2):
@@
Looking at the above equation carefully, it is evident that only one vector field is sufficient to construct the different conserved charges of a black hole. Now we will proceed the same way as we did while deriving the 1st law of BH thermodynamics. We start with the fact that for Killing vectors, the symplectic form vanishes,
@@
So, the flow $v^{a}_{S}$ acts as a conserved current in the thermodynamics of black holes.
First law of cosmological horizons
The first law of thermodynamics for cosmological horizons (in static patch of dS) can also be expressed in terms of bit threads. For de Sitter space, the first law takes the form:
$T_{\rm{dS}}\delta S_{\rm{dS}}=\delta E,$ (4.8)
where $T_{\rm{dS}}=-\frac{1}{2\pi L}$ is the temperature and the minus sign reflects the fact that the cosmological horizon area decreases when energy is added to the system. In terms of perturbative bit threads, this can be written as:
$\nabla_{a}v_{S}^{a}=0,$ (4.9)
similar to the black hole case. This local law have few implications as follows:
The most direct interpretation is that certain concepts related to the thermodynamic quantities associated with the black hole, are locally conserved. This implies that changes in these quantities are not merely due to fluxes across the horizon but are governed by a local balance equation within the bulk itself. This provides a more granular understanding of how black hole (horizon) thermodynamics works.
@@
Here we will discuss how we can rewrite the first law of entanglement entropy in terms of bit threads. Like the previous section, we will use the canonical bit threads to construct a perturbative thread configuration in holographic spacetime. This approach was applied in [17], to reconstruct the metric and to show how linearised Einstein equation can be obtained from perturbative bit threads.
The first law of entanglement entropy says that the change in entanglement entropy is equal to the change in modular energy in first order. In [31], authors showed how one can extend the Iyer-Wald formalism in the subregion duality to reformulate the first law of entanglement entropy in a local way. The change in entanglement entropy and the modular energy can be written in terms of the codim-2 form $\mathbf{k}$ as follows [31]:
$\delta S_{A}=\int_{m(A)}\mathbf{k}_{\xi}\qquad\delta E_{A}=\int_{A}\mathbf{k}_{\xi},$ (4.10)
where, subscript $A$ denotes the boundary subregion for which we are looking at the EE $(S_{A})$, $m(A)$ is the minimal surface homologous to $A$ and $\delta E_{A}$ denotes the change in modular energy. The on-shell condition $\boldsymbol{\omega}=0=\mathrm{d}\mathbf{k}_{\xi}$ leads to the first law $\delta S_{A}=\delta E_{A}$ [^9] This can be easily checked as follows $\int_{\Sigma_{A}}\mathrm{d}\mathbf{k}_{\xi}=0=\int_{\partial\Sigma_{A}}\mathbf{k}_{\xi}\qquad\Rightarrow\int_{m(A)}\mathbf{k}_{\xi}=\int_{A}\mathbf{k}_{\xi}\qquad\Rightarrow\;\delta S_{A}=\delta E_{A}\,.$ . Now using the relation (4.1), we can rewrite the first law of entanglement entropy in terms of the bit thread vector fields,
@@
where $\Sigma_{A}$ denotes the region in between $A$ and $m(A)$, and $\boldsymbol{\epsilon_{\Sigma_{A}}}$ is the volume form of that region. Now, by construction this perturbative bit thread is divergenceless, i.e., the integrand in the above equation does not depend on the size of the region $\Sigma_{A}$. So the first law of the entanglement entropy can be expressed in terms of the bit threads locally as:
@@
Here also thermodynamic flow acts like a conserved current ensuring that the EE and the modular energy are locally conserved.
@@
In a similar way, we can extend the above procedure for differential entropy. But there are few subtleties needed to be considered. Unlike the previous cases, the symplectic current does not vanish, which gives rise to a volume term in the first law of causal diamond [32]. Using the Complexity=Volume``Complexity\,=\,Volume" conjecture, the authors of [33] showed that we can obtain a first law of differential entropy in the holographic scenarios. So the perturbative flow corresponding to the differential entropy will no longer be divergenceless. Unlike the EE and Wald entropy cases, the diffeomorphism generator here is a conformal Killing vector which leads to the non-vanishing of the symplectic current. So, the fundamental identity (2.2) becomes,
@@
where $\chi$ is the conformal Killing vector. The l.h.s. gives rise to the variation of volume of the spatial sphere (hole) whose causal diamond is under consideration and the r.h.s. gives rise to the variation of area of the boundary of that spatial sphere [32]. Integrating on that spatial spherical region $\sigma$,
@@
We can obtain a thermodynamic flow from $\mathbf{k}_{\chi}$, which is a perturbative Riemannian thread. On the other hand we can construct a perturbative Lorentzian thread configuration from the symplectic current using the map: $\delta\mathfrak{u}^{\mu}=g^{\mu\nu}(\ast\boldsymbol{\omega})_{n}$ as discussed in [34], where $\ast$ denotes the Hodge dual in a Lorentzian manifold. In terms of the threads, the above equation becomes,
@@
This gives rise to the first law of differential entropy (or causal diamond). Locally we can write
@@
The above equation gives us a relation between the Lorentzian and Riemannian (perturbative) bit threads.

### v2

@@ 3 Extensions to other entropies
Bit threads can be related to other notions of entropy. In this section, we specifically attempt to relate it to black hole (BH) entropy and differential entropy. Since the CPS formalism is commonly applied in the context of black hole thermodynamics [58, 39], a natural extension is to derive a flow representation for black hole entropy; the same can be done for differential entropy [7, 6], which is defined in both holographic and non-holographic spacetimes.
@@ 3.1 Black hole entropy
The entropy of a black hole has been a topic of interest for a long time, since Bekenstein [9] to the present day. Among various approaches, the definition of black hole entropy as the Noether charge using the CPS formalism, developed by Wald [39, 58], is a well-established and standard notion of entropy, known as the ‘Wald entropy’. We will construct a flow corresponding to the Wald entropy of a black hole. Following the procedure described in section 2.2, we can find a divergenceless vector field. The only difference will be the norm bound at the RT surface. As we are interested in the black hole entropy, the norm bound should saturate on a minimal surface homologous to the black hole horizon (or, more specifically, the bifurcation surface). So the relevant condition becomes $v^{a}=-n_{\mathcal{H}}^{a}/4G_{N}$ at the black hole horizon, $n_{\mathcal{H}}^{a}$ being the unit normal to the horizon. Similar to the AdS-Rindler case, the minus sign emphasises that the direction of $v^{a}$ is towards the horizon. [^11] If we work with a Poincaré like coordinate system, then $v^{a}$ is naturally towards the horizon and we do not need to put the minus sign. Then, the BH entropy is given by,
@@
Here, ‘$M$’ and ‘$J$’ are the mass and angular momentum of the BTZ black hole, respectively; $l$ is the AdS radius. For simplicity, we will consider a non-rotating BTZ black hole, i.e., $J=0$ and $\mathcal{N}^{\phi}=0$. Now, the metric becomes,
@@
where, $f(r)=1-\dfrac{2M}{r}+r^{2}$ with AdS radius $l=1$. Considering spherical symmetries, we can suppress the last term in (3.1). Now the metric becomes,
@@
The rest of the procedure is similar to the last case. Like the previous one, the Schwarzschild-AdS solution is also a solution of Einstein-Hilbert gravity. Then we can use (2.20) to have the codim-2 current. Parametric variation gives us the different components of the variations of the metric as follows,
@@
This $d$-form or the corresponding vector $v^{a}$ related to it via (2.13) gives the black hole entropy upon integration on a codimension-2 surface homologous to the horizon (more specifically, the bifurcation point).
We can think of these flows as the conserved current corresponding to different gravitational charges, and the divergenceless condition is the condition for conservation of those charges, similar to the continuity equation.
@@ 3.2 Differential entropy
In the AdS/CFT correspondence, entanglement entropy in the boundary theory is computed via the Ryu-Takayanagi prescription, which asserts that the entanglement entropy of a boundary region is equal to the minimum area of a surface in the bulk spacetime. This gives us important insights into how quantum information is encoded in spacetime itself. The entanglement entropy extends naturally from individual intervals to overlapping boundary intervals. Each of them has an entanglement entropy given by the Ryu-Takayanagi formula. Taking the continuum limit of differences between these entropies, one introduces differential entropy [6, 35]:
$S_{diff}=\int d\lambda\frac{\partial S\big(\gamma_{L}(\lambda),\gamma_{R}(\lambda)\big)}{\partial\lambda},$ (3.8)
where $\gamma_{L,R}(\lambda)$ are the common endpoints of a family of intervals parameterised by $\lambda$. For particular circumstances, the above integral provides the gravitational entropy of a bulk “hole”. Namely, differential entropy happens to be equal to the area of the hole divided by $4G_{N}$,
@@
Here, we show that the differential entropy corresponding to a hole in a spacetime can be obtained using the flow prescription in spherical Rindler and spherical Rindler-AdS spacetimes. The vector field is derived from CPS formalism with constraints $|v|\leq 1/4G_{N}$, $|v^{a}n_{a}|=1$ at the boundary of the hole, and is divergenceless by definition. So, we can rewrite the differential entropy in terms of the threads as
@@
where $\mathfrak{b}$ is the boundary of the hole and $v=\sqrt{\mathfrak{h_{b}}}\,|n_{a}v^{a}|$. Now, we will look into the explicit construction of the vector field in spherical Rindler space.
@@
The concept of spherical Rindler space generalises the foundational framework of conventional Rindler space, which describes a family of observers undergoing radial acceleration away from a shared central point in Minkowski spacetime, establishing causal isolation from a spherical region of radius $R_{0}$ [7]. Unlike linear Rindler acceleration, which produces planar horizons, this spherical generalisation creates an observationally inaccessible domain bounded by a perfectly symmetrical horizon surface. The spherical Rindler horizon manifests as a topological boundary that separates the causally connected exterior region from the interior zone, which remains perpetually hidden to all accelerated observers. This horizon’s geometrical properties maintain exact spherical symmetry, with its characteristic scale precisely determined by the parameter $R_{0}$, which defines the radius of the inaccessible central region. The $(d+1)$-dimensional metric is given by [7]:
@@
where, $d\Omega_{d-1}^{2}$ denotes the metric on a unit (d1)(d-1)-sphere. This metric structure reveals the intrinsic relationship between the acceleration parameter $R_{0}$, the radial coordinate $r$, and the temporal component that governs the proper time experience of the spherical Rindler observers. The coordinate singularity at $r=R_{0}$ mathematically encodes the causal boundary beyond which no information can propagate to the accelerating observers. The authors of [7] showed that this horizon gives rise to a notion of entropy proportional to the area of the horizon, which was later named as ‘differential entropy’. To obtain the entropy associated with this hole, we generally take the near-horizon limit $r\rightarrow 0$. The metric approaches in the limit
@@
Now we can find a constant time slice in this geometry and employ the Iyer-Wald formalism to find the entropy of the hole, which is essentially given by (3.9). We can also find a flow or bit thread configuration by following the procedure discussed in section 2.2. In the near-horizon geometry, we take the radius of the hole $R_{0}$ as the parameter to be used for the parametric variation of the metric. Using (2.20) and (2.7), we find ,
@@
for the near-horizon geometry of the spherical Rindler space. Here $\mathfrak{c}$ is a function of $R_{0}$. Now using the relations (2.7) and (2.13), we find the vector field,
@@
Notice that the vector field is constant along the $r$ direction, and the minus sign tells that it ends on the horizon of the hole. This reproduces the entropy of the hole as given by (3.9). Therefore, we can see that it is possible to construct a flow-based picture for differential entropy. Moreover, from the above construction, we can say that the CPS formalism can also be applied to obtain the entropy of a hole in a spacetime. Both of these methods are applicable in the near-horizon limit, and the interpretation of the bit thread vector field away from the horizon remains an open question.
@@
Spherical AdS-Rindler space can be viewed as a spherical Rindler space embedded in an asymptotically AdS spacetime. This space is holographically dual to the ultraviolet part of the boundary field theory, described by a finite time strip [6]. This enables us to study a hole in a spacetime holographically. Analogous to the spherical Rindler space, this spacetime also has a family of accelerated observers whose worldlines are causally disconnected from a spherical hole in AdS. The metric of such spacetime is given by [6],
$ds^{2}=l^{2}\left(-\sinh^{2}{r}dt^{2}+dr^{2}\right)+\left(R_{0}\cosh{r}+\sqrt{R_{0}^{2}+l^{2}}\sinh{r}\cosh{t}\right)^{2}d\Omega_{d-1}^{2}\,$, (3.15)
where $R_{0}$ is the coordinate radius of the hole. For convenience, we can take the AdS radius $l$ to be $1$. The horizon is at $r=0$, similar to the spherical Rindler case. Taking the near horizon limit $r\rightarrow 0$, we end up with same metric (3.12):
@@
in the leading order. So, in this geometry, the flow will be the same as given in (3.14). It reproduces the gravitational entropy of the hole, viz., ‘differential’ entropy. This method is only applicable in the near-horizon limit of the hole. If one wants to construct a thread configuration beyond the near-horizon limit, one needs to use the notion of covariant bit threads as formulated in [34].
@@ 4 First laws of entropy using bit-threads
The CPS formalism helps us to construct a representative flow for a valid bit thread configuration. One can wonder if it is possible to represent different thermodynamical laws in the thread picture as well. In this section, we will mainly focus on the first law of thermodynamics for the entropies mentioned above. Before proceeding further, recall that the symplectic current vanishes for Killing symmetries, i.e., $\boldsymbol{\omega}=0=\mathrm{d}\mathbf{k}_{\eta}$. So, we can construct a vector field from the codim-2 current as
@@
which is similar to the map (2.13), and ensures that $v^{a}_{th}$ is divergenceless. We can also impose a norm bound as we did for $v^{a}$ in the previous section. Then we end up with a ‘thermodynamic variation flow’. This flow, when integrated at the boundary of a Cauchy surface, results in variations of different thermodynamic quantities.
@@
For a stationary, rotating, charge-neutral, axially symmetric black hole with non-degenerate Killing vectors and a bifurcation surface, we can derive the first law of black hole thermodynamics using the covariant phase space formalism as briefly discussed in appendix B.3. For such black holes, the Killing vectors associated with the horizon, time translation and rotation are related via the equation (B.21),
@@
where, $\Omega_{H}$ represents the angular velocity of the horizon. Now, it is easy to find out the relation between the codim-2 currents corresponding to the Killing vectors, and it is given by,
@@
Here we have used the notation (A.2). For simplicity, we will omit the suffix ‘$th$’ hereafter. These fields also follow a similar relation like (4.2):
@@
Looking at the above equation carefully, it is evident that only one vector field is sufficient to construct the different conserved charges of a black hole. Now we will proceed the same way as we did while deriving the first law of BH thermodynamics. We start with the fact that for Killing vectors, the symplectic form vanishes,
@@
So, the flow $v^{a}_{S}$ acts as a conserved current in the thermodynamics of black holes. The most direct interpretation of this local law is that certain concepts related to the thermodynamic quantities associated with the black hole are locally conserved. It implies that changes in these quantities are not merely due to fluxes across the horizon but are governed by a local balance equation within the bulk itself. This provides a more granular understanding of how black hole thermodynamics works.
@@
Here, we will discuss how we can rewrite the first law of entanglement entropy in terms of bit threads. Similar to the previous section, we will use the canonical bit threads to construct a perturbative thread configuration in holographic spacetime. This approach was applied in [2] to reconstruct the metric and to show how linearised Einstein equations can be obtained from perturbative bit threads.
The first law of entanglement entropy says that the change in entanglement entropy is equal to the change in modular energy in the first order. In [26], authors showed how one can extend the Iyer-Wald formalism in the subregion duality to reformulate the first law of entanglement entropy in a local way. The change in entanglement entropy and the modular energy can be written in terms of the codim-2 form $\mathbf{k}$ as follows [26]:
$\delta S_{A}=\int_{m(A)}\mathbf{k}_{\xi}\;;\quad\delta E_{A}=\int_{A}\mathbf{k}_{\xi},$ (4.8)
where, subscript $A$ denotes the boundary subregion for which we are looking for the EE, $m(A)$ is the minimal surface homologous to $A$ and $\delta E_{A}$ denotes the change in modular energy. The on-shell condition $\boldsymbol{\omega}=0=\mathrm{d}\mathbf{k}_{\xi}$ leads to the first law $\delta S_{A}=\delta E_{A}$. [^12] It can be easily checked as follows $\int_{\Sigma_{A}}\mathrm{d}\mathbf{k}_{\xi}=0=\int_{\partial\Sigma_{A}}\mathbf{k}_{\xi}\qquad\Rightarrow\int_{m(A)}\mathbf{k}_{\xi}=\int_{A}\mathbf{k}_{\xi}\qquad\Rightarrow\;\delta S_{A}=\delta E_{A}\,.$ Now using the relation (4.1), we can rewrite the first law of entanglement entropy in terms of the bit thread vector fields,
@@
where $\Sigma_{A}$ denotes the region in between $A$ and $m(A)$, and $\boldsymbol{\epsilon_{\Sigma_{A}}}$ is the volume form of that region. Now, by construction, this perturbative bit thread is divergenceless, i.e., the integrand in the above equation does not depend on the size of the region $\Sigma_{A}$. So the first law of the entanglement entropy can be expressed in terms of the bit threads locally as:
@@
Here, thermodynamic flow acts like a conserved current, ensuring that the EE and the modular energy are locally conserved.
@@
In a similar way, we can extend the above procedure for differential entropy. But there are a few subtleties that need to be considered. Unlike the previous cases, the symplectic current does not vanish and gives rise to a volume term in the first law of causal diamond [40]. Using the Complexity=Volume``Complexity\,=\,Volume" conjecture, the authors of [54] showed that one can obtain a first law of differential entropy in the holographic scenarios. So the perturbative flow corresponding to the differential entropy will no longer be divergenceless. Unlike the EE and Wald entropy cases, the diffeomorphism generator here is a conformal Killing vector, which leads to the non-vanishing of the symplectic current. So, the fundamental identity (2.2) becomes,
@@
where $\chi$ is the conformal Killing vector. The l.h.s. gives rise to the variation of volume of the spatial sphere (hole) whose causal diamond is under consideration, and the r.h.s. gives rise to the variation of the area of the boundary of that spatial sphere [40]. Integrating over that spatial spherical region $\sigma$,
@@
We can obtain a thermodynamic flow from $\mathbf{k}_{\chi}$, which is a perturbative Riemannian thread. On the other hand, we can construct a perturbative Lorentzian thread configuration from the symplectic current using the map: $\delta\mathfrak{u}^{\mu}=g^{\mu\nu}(\ast\boldsymbol{\omega})_{n}$ as discussed in [49], where $\ast$ denotes the Hodge dual in a Lorentzian manifold. In terms of the threads, the above equation becomes,
@@
This gives rise to the first law of differential entropy (or causal diamond). Locally, we can write
@@
The above equation provides a relation between the Lorentzian and Riemannian (perturbative) bit threads.